OpenStudy (anonymous):

Line that is parallel to the like 2y-4x=10 and goes through the point (-3,-7)

OpenStudy (anonymous):

Lines that are parallel have the same slope so first you must put the equation into y=mx+b (slope intercept form) \[2y=4x+10\] \[y=2x+5\] now that you know the slope (m) , you can use the point slope form \[y-y_1=m(x-x_1)\] where \[(x_1,y_1)\] is a point on the line you are trying to find the equation to

OpenStudy (anonymous):

yes i need to find the equation

OpenStudy (anonymous):

i got y=2x+1 is this right?

OpenStudy (anonymous):

\[y-(-7)=2(x-(-3)\] \[y+7=2(x+3)\] \[y+7=2x+6\] \[y+7-7=2x+6-7\] \[y=2x-1\]

OpenStudy (anonymous):

thanks